A new concrete water tank (unit weight= 24 kN/m3) with inner diameter of 13.8 m with 0.1 m wall thickness, and 9.0 m high is proposed for construction to cover the water shortage of government officers’ quarters of UNIMAS that located near to an existing 6.0 m high concrete water tank with similar diameter. The existing water tank sits on a concrete foundation of diameter 15.0 m and thickness of 0.6 m which is below the finished surface level. The existing water tank has fully loaded to its capacity up to a height 5.8 m (unit weight = 9.81 kN/m3) during the construction of the new tank. The new water tank is located at a distance of 4.0 m centre-centre to the existing concrete tank. The new tank sits on a similar foundation as the former. The foundation bottom is 0.6 below the finished surface elevation. The soil consists of 5m thick soft, overconsolidated silty clay and overlain by 1.0 m silty sand. The clay layer sits above a very thick layer of completely weathered mudstone, more than 10 m thick. The groundwater level is 0.6 m below the surface level. The geotechnical data of the soft clay gave Cc = 0.28, Cr = 0.06, Cv = 7.427 × 10−5 m2 / day, wclay = 40%, wsand = 30% and OCR = 8. The specific gravity of sand and clay are 2.6 and 2.7, respectively. Assume Poisson ratio, v = 0.35. Note: if at any point in this examination you feel that anything is unclear, please make any additional reasonable assumptions that you feel necessary and please state it clearly. State any additional sources i.e literature you have used in making assumptions. a) Estimate the immediate settlement under the centre of new tank just after the construction with both tanks full. b) After 6 months of service having both tanks are full, estimate the primary consolidation settlement of the clay layer under the centre of new tank. c) As an engineer, comment on the impact of the soil total settlement due to both tanks considering allowable settlement is 25 mm in this construction.
Answer #1

(a) Diagnam ces per given data : - water tank 13.8m Im 14m- 0-1 m Thickness of wall G.L. 9 Di 0-6m ܚ ✓ Y = 24kw/m3 • ConcreteWe know that weight of toundation - Id2x0.6x24 x 152x0•6X24 = ㅈ 4 = 2543.4kn weight of tank = * 4 (142-13.92)x9x24 Kolm? 942-Elastic settlement of silt layer:- We know that pe = AG XD (1-72) X Ich Esec +94.46X14 (1-0.35?) x 1 933665 olm? (Assume) = 0Ar=qsIc = 94.46800724 68040 KN/m2 pez ARD (1-02) 38304:20 다. 168.40x14 (1-0-352) 38304.20 Pe=0.021m (b) Long term for consoliA62 = 94-46 X 0.676=63094 kN]m2 We know that oftm = 6żotar =40+ 63.94 = 103.94 kN/m² <ozó Ppc = Ho Iteo teor log (log oftm =